### v2

# hyperrefIgnoring empty anchor

Abstract
This paper considers a novel variant of the online fair division problem involving multiple agents in which a learner sequentially observes an indivisible item that has to be irrevocably allocated to one of the agents while satisfying a fairness and efficiency constraint. Existing algorithms assume a small number of items with a sufficiently large number of copies, which ensures a good utility estimation for all item-agent pairs from noisy bandit feedback. However, this assumption may not hold in many real-life applications, for example, an online platform that has a large number of users (items) who use the platform’s service providers (agents) only a few times (a few copies of items), which makes it difficult to accurately estimate utilities for all item-agent pairs. To address this, we assume utility is an unknown function of item-agent features. We then propose algorithms that model online fair division as a contextual bandit problem, with sub-linear regret guarantees. Our experimental results further validate the effectiveness of the proposed algorithms.
1 Introduction
Growing economic, environmental, and social pressures require us to be efficient with limited resources (Aleksandrov and Walsh, 2020). Therefore, the fair division (Steinhaus, 1948) of limited resources among multiple parties/agents is needed to efficiently balance their competing interests in
Figure 1: Example of an online fair division of numerous items with few copies: An online platform recommends a service provider (agent) to sequentially arriving users (items). The platform must balance two conflicting objectives: fairly recommending service providers to address their competing interests (fairness) and maximizing its own profit (efficiency).
many real-life applications, for example, Fisher market (Codenotti and Varadarajan, 2007; Vazirani, 2007), housing allocation (Benabbou et al., 2019), rent division (Edward Su, 1999; Gal et al., 2016), and many more (Demko and Hill, 1988). The fair division problem has been extensively studied in algorithmic game theory (Eisenberg and Gale, 1959; Codenotti and Varadarajan, 2007; Vazirani, 2007; Caragiannis et al., 2019) but focuses on the static setting where all information (items, agents, and their utility) is known in advance. However, many real-life fair division problems are often online (Aleksandrov and Walsh, 2020; Gao et al., 2021; Gkatzelis et al., 2021; Benadè et al., 2022; Liao et al., 2022; Yamada et al., 2024; Yang et al., 2024), referred to as online fair division, where indivisible items arrive sequentially and each item needs to be irrevocably allocated to an agent. Existing algorithms for online fair division assume a small number of items with a sufficiently large number of copies (Yamada et al., 2024), ensuring a good utility estimation using the observed utilities for previous allocations for all item-agent pairs. These estimated utilities are used to find the agent for allocating the item that maintains a desired balance between fairness (i.e., keeping the desired level of utilities across the agents) and efficiency (i.e., maximizing the total utility) (Sinclair et al., 2022).
However, many real-life applications have a large number of items with only a few copies for each item. For example, as illustrated in Fig. 1, consider an online food delivery platform that wants to recommend restaurants (agents) to its users (items) while balancing between fairly recommending restaurants to accommodate their competing interests (fairness) and maximizing its own profit (efficiency). Similar scenarios also arise while recommending the cab to users by ride-hailing platforms, e-commerce platforms choosing top sellers to buyers, network resource allocation (Lan et al., 2010; Verma and Hanawal, 2020), online advertisement (Li et al., 2019), allocating tasks among crowd-sourced workers (Patil et al., 2021; Yamada et al., 2024), providing humanitarian aid post-natural disaster (Yamada et al., 2024), among many others (Walsh, 2011; Mehta et al., 2013; Aleksandrov and Walsh, 2020). Having a large number of items with only a few copies of each item makes it impossible to estimate the utility for all item-agent pairs. Therefore, this problem raises a natural question: How to design an algorithm for online fair division problems having a large number of items with only a few copies for each item?
The closest works to ours are (Bhattacharya et al., 2024; Procaccia et al., 2024; Yamada et al., 2024; Schiffer and Zhang, 2025), but these works assume that the online fair division problem has a small number of items and agents with a sufficiently large number of copies of each item to ensure a good utility estimation for all item-agent pairs. In contrast, we consider an online fair division setting where the number of items can be large while having only a few copies for each item, thus considering a more general problem setting than in existing works, making existing algorithms (Bhattacharya et al., 2024; Procaccia et al., 2024; Yamada et al., 2024; Schiffer and Zhang, 2025) ineffective in our setting, as accurately estimating utility for all item-agent pairs is difficult or even infeasible. Therefore, the first challenge we faced in designing an algorithm for our setting is how to estimate the utility for any item-agent pair using observed stochastic utilities from previous allocations so that these estimated utilities can be used to find the best agent for allocating an item observed in the subsequent round. A natural solution is to assume a correlation structure exists between the item-agent features and their utility, which can be realized via parameterization of the expected utility such that the utility depends on the unknown function of item-agent features.
Next challenge we faced is that, in many real-life applications, agents may also be concerned about fairness in each round of the allocation process (Sim et al., 2021; Benadè et al., 2023), for example, in the online platform, if service providers (agents) do not get an adequate number of users (items), they may choose to leave the platform, or even worse, may join another competing platform. Being fair in each round leads to the next challenge of finding how well the item’s allocation to an agent will maintain the desired balance between fairness and efficiency. To address this challenge, we introduce the notion of goodness function that measures how well an item allocation to an agent maintains the desired balance between fairness and efficiency (a larger goodness function value implies a better allocation). Thus, the algorithm’s goal is to allocate each item to the agent that maximizes the goodness function value.
Even when equipped with an estimate of the unknown utility function and goodness function, the algorithm must decide which agent should be allocated the given item. Since the utility is only observed for the agent to whom item is allocated (i.e., item-agent pair), the algorithm has to deal with the exploration-exploitation trade-off (Auer et al., 2002; Slivkins, 2019; Lattimore and Szepesvári, 2020), i.e., choosing the best agent based on the current utility estimate (exploitation) or trying a new agent that may lead to a better utility estimates in the future (exploration). To deal with the exploration-exploitation trade-off, we adapt the contextual bandit algorithms (Li et al., 2010; Chu et al., 2011; Krause and Ong, 2011; Agrawal and Goyal, 2013; Chowdhury and Gopalan, 2017; Li et al., 2017; Verma et al., 2020; Zhou et al., 2020; Zhang et al., 2021; Verma et al., 2023, 2025) to get an optimistic estimate of unknown utility for each item-agent pair (in Section 4), where the reward in contextual bandits corresponds to utility, contexts to items, and arms to agents. The goodness function (for example, weighted Gini social-evaluation function (Weymark, 1981) or locally monotonically non-decreasing and Lipschitz functions, more details are in Section 3 and Section 4.3) uses these optimistic utility estimates to allocate the given item to the best possible agent that maintains the desired balance between fairness and efficiency. To measure the algorithm’s performance in our setting, we use the notion of regret, which is the sum of the difference between the maximum goodness function value (i.e., selecting the optimal agent) and goodness function value after the algorithm selects an agent for an item. Specifically, our key contributions are as follows:
  • We introduce the online fair division problem with noisy bandit feedback in which there are a large number of items while each item only having a few copies. We then introduce the notion of goodness function, then define regret for our problem setting in Section 2, and finally propose algorithms with sub-linear regret guarantees in Section 4.
  • We first prove that our proposed algorithms achieve a sub-linear regret upper bound when the goodness function is the weighted Gini social evaluation function (1 and 2), then extend this result to a class of goodness functions that are locally monotonically non-decreasing and locally Lipschitz continuous (3 in Section 4.3).
  • Our experimental results in Section 5 also corroborate our theoretical results and validate the different performance aspects of our proposed algorithms.
1.1 Related Work
In this section, we focus on the most relevant work on online fair division and cover related topics, such as fair division and fair multi-armed bandits, which are discussed in the supplementary material.
Online fair division.  The online fair division involves problems in which items arrive sequentially and then must be immediately and irrevocably allocated to one of the agents. The nature of items can be different: either divisible (Walsh, 2011) or indivisible (Aleksandrov and Walsh, 2017; Procaccia et al., 2024; Yamada et al., 2024), homogeneous/heterogeneous (Walsh, 2011; Kash et al., 2014), multiple copies of items (Procaccia et al., 2024; Yamada et al., 2024), and agents receiving more than one item (Aleksandrov et al., 2015). However, finding a fair allocation is also closely related to computing market or competitive equilibria, both of which are known to be computationally challenging intractable in general (Codenotti and Varadarajan, 2007; Vazirani, 2007). In a static setting with all items available upfront, this problem can be formulated as the Eisenberg-Gale (EG) convex program (Eisenberg and Gale, 1959; Jain and Vazirani, 2010). In online settings, envy-freeness (EF) is incompatible with Pareto optimality (PO), even with divisible items. Some works relax EF with approximate solutions, such as relaxed EF (Kash et al., 2014), stochastic approximation for the EG program (Bateni et al., 2022), or approximate fairness (Yamada et al., 2024).
Fairness and efficiency.  Recent works (Sinclair et al., 2022; Benadè et al., 2023) have examined the trade-off between fairness and efficiency in online allocation problems. Specifically, prior work (Benadè et al., 2023) assumes that the utilities of all agents for an item are known upon its arrival (noiseless utility). In the online fair allocation problem with partial information, where the algorithm observes ordinal rankings rather than cardinal values, it has been shown that EF and approximate PO can coexist (Benadè et al., 2022). In contrast, we consider a setting where utilities are unknown and only noisy feedback is observed for an item allocation, leading to a learning problem. To address this, existing works (Bhattacharya et al., 2024; Procaccia et al., 2024; Yamada et al., 2024; Schiffer and Zhang, 2025) model online fair division problems as a multi-armed bandits problem, but only consider problems having a small number of items and agent with a sufficiently large number of copies of each item to ensure a good utility estimation for all item-agent pairs. Whereas, we consider a more general setting where the number of items can be large but only a few copies of each item are available, making existing algorithms ineffective.
2 Problem Setting
Online fair division.  We consider an online fair division problem involving $N$ agents in which a learner (or central planner) only observes an indivisible item each round that must allocated to one of the agents while satisfying a given fairness constraint to the greatest extent possible. We denote the set of agents by $\mathcal{N}$ and the set of indivisible items by $\mathcal{M}$. In our problem, we can have a large number of items with only a few copies of each item, and the learner has no prior knowledge about future items. At the start of the round $t$, the environment selects an item $m_{t}\in\mathcal{M}$ (which is drawn from an unknown distribution $\nu$), and then the learner observes and allocates the item $m_{t}$ to an agent $n_{t}\in\mathcal{N}$. After that allocation, the learner observes a stochastic utility collected by the agent, denoted by $y_{t}\doteq f(m_{t},n_{t})+\varepsilon_{t}$, where $y_{t}\in\mathbb{R}^{+}$, $f:\mathcal{M}\times\mathcal{N}\rightarrow\mathbb{R}^{+}$ is an unknown utility function, and $\varepsilon_{t}$ is a $R$-sub-Gaussian noise, i.e., $\forall\lambda\in\mathbb{R},~{}~{}~{}\mathbb{E}\left[e^{\lambda\varepsilon_{t}}|\left\{m_{s},n_{s},\varepsilon_{s}\right\}_{s=1}^{t-1},m_{t},n_{t}\right]\leq\exp\left({\lambda^{2}R^{2}}/{2}\right)$.
Allocation quality measure.  Let $U_{t}^{n}$ be the cumulative total utility collected by agent $n$ at the beginning of the round $t$, where $U_{t}^{n}\doteq\sum_{s=1}^{t-1}y_{s}\mathbbmss{1}{\left(n_{s}=n\right)}$ and $\mathbbmss{1}{\left(n_{s}=n\right)}$ denotes the indicator function. We denote the vector of all agents’ total utility by $\bm{U}_{t}$, where $\bm{U}_{t}\doteq\left(U_{t}^{n}\right)_{n\in\mathcal{N}}.$ We assume a goodness function G exists that incorporates the desired level needed between fairness and efficiency. This goodness function measures how well the item allocation to an agent will maintain the desired balance between fairness and efficiency (larger value implies better allocation). We discuss the different choices of the goodness function in Section E.1. For the given total utilities of agents, the value of the goodness function G after allocating the item $m_{t}$ to an agent $n$ is denoted by $\textrm{G}\left(\bm{U}_{t,n}\right)$. Here, $\bm{U}_{t,n}$ is the same as $\bm{U}_{t}$ except the total utility of $n$-th agent is $U_{t}^{n}+f(m_{t},n)$. The learner’s goal is to allocate the item to an agent that maximizes the value of the goodness function or satisfies the given fairness constraint to the greatest extent possible.
Performance measure of allocation policy.  Let $n_{t}^{\star}$ denote the optimal agent for item $m_{t}$ having the maximum goodness function value, i.e., $n_{t}^{\star}=\arg\!\max_{n\in\mathcal{N}}\left[\textrm{G}\left(\bm{U}_{t,n}\right)\right]$. Since the utility function $f$ is unknown, we cannot directly compute the optimal agent $n_{t}^{\star}$ for item $m_{t}$. To overcome this, we sequentially estimate the utility function $f$ using the historical information of the stochastic utility observed for the item-agent pairs and then use the estimated utility function to allocate an agent $(n_{t})$ for the item $m_{t}$. After allocating item to an agent $n_{t}$, the learner incurs a penalty (or instantaneous regret) $r_{t}$, where $r_{t}=\textrm{G}\left(\bm{U}_{t,n_{t}^{\star}}\right)-\textrm{G}\left(\bm{U}_{t,n_{t}}\right).$ We use this notion of penalty as a performance measure because our goal is to achieve a fair allocation by the end of each round, which is motivated by practical applications where the learner must satisfy a desired fairness and efficiency constraint in each item allocation. Our aim is to learn a sequential policy that selects agents for items such that the learner’s total penalty for not assigning the item to the optimal agent (or cumulative regret) is as minimal as possible. Specifically, the cumulative regret (regret in short for brevity) of a sequential policy $\pi$ that selects agent $n_{t}$ for allocating item $m_{t}$ in the round $t$ in the $T$ rounds is given by
$\mathfrak{R}_{T}(\pi)\doteq\sum_{t=1}^{T}r_{t}=\sum_{t=1}^{T}\left[\textrm{G}\left(\bm{U}_{t,n_{t}^{\star}}\right)-\textrm{G}\left(\bm{U}_{t,n_{t}}\right)\right].$ (1)
A policy $\pi$ is considered a good policy if it has sub-linear regret, i.e., $\lim_{T\rightarrow\infty}{\mathfrak{R}_{T}(\pi)}/T=0$. This implies that the policy $\pi$ will eventually start allocating items to the optimal agent. Note that our regret definition differs from standard contextual bandits algorithms, as the value of the goodness function depends not only on the expected utility in the current round (like in contextual bandits) but also on the utilities collected by each agent in the past (history-dependent). This dependence on history makes regret analysis challenging for any arbitrary goodness function. We provide a detailed discussion of the above regret definition and its different aspects in Appendix B.
3 Goodness Function Incorporating Fairness and Efficiency Constraint
In the following, we define the various performance measures for fairness and efficiency commonly used in the algorithmic game theory and economics. For brevity, let $U_{n}$ be the utility of agent $n$.
  • $\blacksquare$
    Utilitarian Social Welfare (Feldman and Serrano, 2006) . Utilitarian Social Welfare (USW) is defined as the sum of utilities of all agents, i.e., $\sum_{n\in\mathcal{N}}U_{n}$. Maximizing USW indicates the system’s total utility is maximum. We refer to this as an efficiency measure.
  • $\blacksquare$
    Egalitarian Social Welfare (Feldman and Serrano, 2006). Egalitarian Social Welfare (ESW) is defined as the minimum utility across all agents, i.e., $\min_{n\in\mathcal{N}}U_{n}$. It is a notion of fairness where the designer aims to maximize the utility of less happy agents to obtain a fair outcome.
  • $\blacksquare$
    Nash Social Welfare (Nash et al., 1950). Nash Social Welfare (NSW) is defined as the geometric mean of the utilities of all agents, i.e., $\left[\prod_{n\in\mathcal{N}}U_{n}\right]^{{1}/{|\mathcal{N}|}}$. By maximizing NSW, we obtain an approximately fair and efficient allocation that achieves a balance between USW and ESW.
Goodness function.  Since the goodness function is a performance indicator of the item allocation to an agent, it is used to allocate the items to agents in a way that maintains the desired balance between fairness and efficiency constraint. It is a well-known fact that there is a trade-off between these two performance measures. For instance, when a learner aims to maximize social welfare, the item is allocated to an agent with the highest utility. Such an allocation scheme achieves efficiency but sacrifices fairness. Intuitively, one way to obtain fairness is to apply a smaller weight factor to agents with higher cumulative utility and a larger weight factor to those with lower cumulative utility. Therefore, we must consider an appropriate goodness function G such that optimizing G corresponds to a) maximizing efficiency, i.e., maximizing the individual agent’s cumulative utility, and b) reducing the utility disparity among the agents, which ensures fairness. Our regret, as defined in Eq. 1, is formulated for a general goodness function G. We first only consider the goodness function to be the weighted Gini social-evaluation function (Weymark, 1981), which measures the weighted fairer distribution of utility while balancing the trade-off between fairness and efficiency. To define this function, we introduce a positive and non-increasing weight vector $\bm{w}_{\mathcal{N}}=(w_{1},\ldots,w_{|\mathcal{N}|})$, where $0\leq w_{n}\leq 1$, and $\Phi_{n}$ that is a sorting operator and arranges the cumulative utilities in increasing order. The weighted Gini social-evaluation function is then given as follows:
$\textrm{G}\left(\bm{U}_{t,n_{t}}\right)=\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n_{t}}\right).$ (2)
We now consider the following cases that coincide with the well-known social welfare functions.
  1. 1.
    If $w_{1}=1$ and $w_{n}=0$ for $n\geq 2$, then the weighted Gini social-evaluation function aligns with ESW, i.e., maximizing the minimum utility among agents.
  2. 2.
    If $w_{n}=1$ for all $\forall n\in\mathcal{N}$, then the weighted Gini social-evaluation function aligns with USW, i.e., maximizing the goodness function promotes efficiency.
  3. 3.
    If $0<w_{n}\leq 1$ for all $n\in\mathcal{N}$, then by appropriately choosing the weights, we can effectively control the trade-off between efficiency and fairness.
Remark 1.
In the first two cases, we observe that larger variability in the weights leads to fairness, while smaller variability promotes efficiency. To control the trade-off with a single parameter instead of $|\mathcal{N}|$ parameters, we can set the weight $w_{n}=\rho^{n-1}$, for all $n\in\mathcal{N}$, where $0<\rho\leq 1$ is a control parameter. We experimentally study the impact of $\rho$ on fairness and efficiency in Section 5.
We primarily focus on the weighted Gini social-evaluation function because it effectively allows us to interpolate between fairness (ESW) and efficiency (USW) by adjusting a single parameter $\rho$, as discussed in 1. We also extend our results to other classes of goodness functions, such as the NSW and log-NSW; further details are provided in Section 4.3 and Section E.1.
4 Contextual Bandits for Online Fair Division
The online fair division setting that we consider can have a large number of items with only a few copies (even only one); hence, it is impossible to get a good utility estimate for all item-agent pairs. To overcome this challenge, we assume a correlation between the utility and item-agent features, which can be realized via parameterization of the expected utility such that the utility depends on the unknown function of item-agent features. This assumption is common in the contextual bandit literature (Li et al., 2010; Chu et al., 2011; Agrawal and Goyal, 2013; Zhou et al., 2020; Zhang et al., 2021), where the reward (utility) is assumed to be an unknown function of context-arm (item-agent) features. In this paper, we model online fair division of a large number of items with a few copies as a contextual bandit problem in which the item is treated as the context, the agent as the arm, and utility as the reward. To bring out our key ideas and results, we first focus on linear utility functions and later extend our results to non-linear functions in Section 4.2.
4.1 Linear Utility Function
For brevity, let $\mathcal{M}_{t}\subset\mathbb{R}^{d}$ $(d\geq 1)$ denote the set of all item-agent feature vectors in the round $t$, and $m_{t,n}$ represent the item-agent features for item $m_{t}$ and an agent $n$. After allocating the item $m_{t}$ to an agent $n_{t}$, the learner observes stochastic utility $y_{n_{t}}=m_{t,n_{t}}^{\top}\theta^{\star}+\varepsilon_{t}$, where $\theta^{\star}\in\mathbb{R}^{d}$ is the unknown parameter and $\varepsilon_{t}$ is $R$-sub-Gaussian. Let $M_{t}\doteq\sum_{s=1}^{t-1}m_{s,n_{s}}m_{s,n_{s}}^{\top}+\lambda I_{d}$, where $\lambda>0$ is the regularization parameter that ensures $M_{t}$ is a positive definite matrix and $I_{d}$ is the $d\times d$ identity matrix. The weighted $l_{2}$-norm of vector $m$ with respect to matrix $M$ is denoted by $\left\|m\right\|_{M}$.
At the start of round $t$, $\hat{\theta}_{t}={M}_{t}^{-1}\sum_{s=1}^{t-1}m_{s,n_{s}}y_{n_{s}}$ is the estimate of the unknown parameter $\theta^{\star}$. After having this utility function estimator, the learner has to decide which agent to allocate the given item. Since the utility is only observed for the selected item-agent pair, the learner needs to deal with the exploration-exploitation trade-off (Auer et al., 2002; Lattimore and Szepesvári, 2020), i.e., choosing the best agent based on the current utility estimate (exploitation) or trying a new agent that may lead to a better utility estimator in the future (exploration). The upper confidence bound (UCB) (Li et al., 2010; Chu et al., 2011; Zhou et al., 2020) and Thompson sampling (TS) (Krause and Ong, 2011; Agrawal and Goyal, 2013; Zhang et al., 2021) are widely-used techniques for dealing with the exploration-exploitation trade-off.
OFD-UCB UCB-based algorithm for online fair division with linear utility function
1:  Input: $\lambda>0$, Number of agents $N=|\mathcal{N}|$
2:  For the first $N$ rounds: allocates items to agents in a round-robin fashion
3:  Compute $M_{N+1}=\lambda I_{d}+\sum_{s=1}^{N}m_{s,n_{s}}m_{s,n_{s}}^{\top}$ and $\hat{\theta}_{N+1}={M}_{N+1}^{-1}\sum_{s=1}^{N}m_{s,n_{s}}y_{n_{s}}$
4:  for $t=N+1,N+2,\ldots$ do
5:     Observe an item $m_{t}$
6:     Select agent $n_{t}\in\arg\!\max_{n\in\mathcal{N}}\textrm{G}\left(\bm{U}_{t,n}^{\text{UCB}}\right)$. Break ties randomly
7:     Observe stochastic utility $y_{n_{t}}$ for agent $n_{t}$
8:     Update $M_{t+1}=M_{t}+m_{t,n_{t}}m_{t,n_{t}}^{\top}$ and re-estimate $\hat{\theta}_{t+1}={M}_{t+1}^{-1}\sum_{s=1}^{t}m_{s,n_{s}}y_{n_{s}}$
9:  end for
UCB-based algorithm.  We first propose a UCB-based algorithm, named OFD-UCB, that works as follows. In the first $N$ rounds, the learner allocates items to agents in a round-robin fashion to ensure each agent has positive utility. At round $t$, the environment reveals an item $m_{t}$ to the learner. Before selecting the agent for that item, the learner updates the utility function estimate $(\hat{\theta}_{t})$ using available historical information before the round $t$ (i.e., $\{m_{s,n_{s}},y_{n_{s}}\}_{s=1}^{t-1}$). Then, the utility UCB value for allocating item $m_{t}$ to an agent $n$, denoted by $u_{m_{t},n}^{\text{UCB}}$, is computed as follows:
$u_{m_{t},n}^{\text{UCB}}=m_{t,n}^{\top}\hat{\theta}_{t}+\alpha_{t}\left\|m_{t,n}\right\|_{{M}_{t}^{-1}},$ (3)
where $m_{t,n}^{\top}\hat{\theta}_{t}$ is the estimated utility for allocating item $m_{t}$ to agent $n$ and $\alpha_{t}\left\|m_{t,n}\right\|_{{M}_{t}^{-1}}$ is the confidence bonus in which $\alpha_{t}=R\sqrt{d\log\left(\nicefrac{{1+\left({tL^{2}}/{\lambda}\right)}}{{\delta}}\right)}+\lambda^{\frac{1}{2}}S$ is a slowly increasing function in $t$ and the value of $\left\|m_{t,n}\right\|_{{M}_{t}^{-1}}$ goes to zero as $t$ increases.
Using this optimistic utility estimate of each agent, the algorithm selects an agent for allocating the item $m_{t}$ by maximizing the optimistic value of the goodness function as follows:
$n_{t}\in\arg\!\max_{n\in\mathcal{N}}\textrm{G}\left(\bm{U}_{t,n}^{\text{UCB}}\right),$ (4)
where $\bm{U}_{t,n}^{\text{UCB}}=\left(U_{t}^{a}+u_{m_{t},a}^{\text{UCB}}\mathbbmss{1}{\left(a=n\right)}\right)_{a\in\mathcal{N}}$ in which $U_{t}^{a}$ is the total utility collected by the agent $a$ before the round $t$. If there are multiple agents to whom allocating the item gives the same maximum value of goodness function, then one of these agents is chosen randomly. After allocating item $m_{t}$ to agent $n_{t}$, the environment generates a stochastic utility $y_{n_{t}}$. The learner observes the utility $y_{n_{t}}$ and then updates the values of $M_{t+1}=M_{t}+m_{t,n_{t}}m_{t,n_{t}}^{\top}$ and re-estimates $\hat{\theta}_{t+1}={M}_{t+1}^{-1}\sum_{s=1}^{t}m_{s,n_{s}}y_{n_{s}}$. The same process is repeated to select agents for subsequent items. Our next result gives the regret upper bound of OFD-UCB when the utility function is assumed linear.
Theorem 1.
Let $\delta\in(0,1)$, $\lambda>0$, noise in utility be the $R$-sub-Gaussian, and the goodness function be same as defined in Eq. 2 with $w_{\max}=\max_{n\in\mathcal{N}}{w_{n}}$. Then, with a probability of at least $1-\delta$, the regret in $T>0$ rounds is
$\mathfrak{R}_{T}\left(\textnormal{{\bf OFD-UCB}{}}\right)\leq 2\alpha_{T}w_{\max}\sqrt{2dT\log(\lambda+TL/d)},$
where $\alpha_{T}=R\sqrt{d\log\left(\frac{1+\left({TL^{2}}/{\lambda}\right)}{\delta}\right)}+\lambda^{\frac{1}{2}}S$, $\left\|\theta^{\star}\right\|_{2}\leq S$, and $\left\|m_{t,n}\right\|_{2}\leq L~{}\forall t\geq 1,n\in\mathcal{N}$.
Proof outline.
The key observation is cumulative regret depends on the instantaneous regret incurred in each round. We can upper bound the instantaneous regret $(r_{t})$ incurred in the round $t$ by $2w_{\max}\left\|\hat{\theta}_{t}-\theta^{\star}\right\|_{{M}_{t}}\left\|x_{t}\right\|_{{M}_{t}^{-1}}$. As the instantaneous regret depends on the estimation error of $\theta^{\star}$ using observed item-agent pairs, i.e., $\left\|\hat{\theta}_{t}-\theta^{\star}\right\|_{{M}_{t}}$ in the round $t$, we adapted results from linear contextual bandits in our setting to get an upper bound on this estimation error. With this result, the regret upper bound follows by upper-bounding the sum of all instantaneous regret. The detailed proofs of 1 and proofs of other related results are provided in Appendix C.
TS-based algorithm.  Due to the empirical superiority of TS-based algorithms over UCB-based bandit algorithms (Chapelle and Li, 2011; Agrawal and Goyal, 2013), we propose a TS-based algorithm named OFD-TS, which works similar to OFD-UCB except the agent selection (Line 6). To get a TS-based utility estimate, the algorithm first samples a utility function parameter $\tilde{\theta}_{t}$ from the distribution $\mathcal{N}\left(\hat{\theta}_{t},\beta_{t}^{2}V_{t}^{-1}\right)$, where $\mathcal{N}$ denotes the normal distribution and $\beta_{t}=R\sqrt{9d\log\left(t/\delta\right)}$ (Agrawal and Goyal, 2013). Using $\tilde{\theta}_{t}$, the TS-based utility estimate, i.e., $u_{m_{t},n}^{\text{TS}}=m_{t,n}^{\top}\tilde{\theta}_{t}$, replaces $u_{m_{t},n}^{\text{UCB}}$ for computing the value of goodness function in Eq. 4.
4.2 Non-linear Utility Function
We now consider the setting in which the utility function can be non-linear. As shown in  Section 4.1, the linear contextual bandit algorithm can be used as a sub-routine to get the optimistic utility estimate in our contextual online fair division problem. These estimates are then used to compute the value of the goodness function for allocating the observed item to the agent that gives the best-desired balance between fairness and efficiency. We generalize this observation for the online fair division problem with a non-linear utility function by using a suitable non-linear contextual bandit algorithm and introduce the notion of Online Fair Division (OFD) Compatible contextual bandit algorithm.
Definition 1 (OFD Compatible Contextual Bandit Algorithm).
Let $\mathcal{O}_{t}$ denote the observations of item-agent pairs at the beginning of round $t$ and $m_{t,n}\in\mathcal{M}$. Then, any contextual bandit algorithm $\mathfrak{A}$ is OFD Compatible if its estimated function $f_{t}^{\mathfrak{A}}$, with probability $1-\delta$, satisfies:
$|f_{t}^{\mathfrak{A}}(m_{t,n})-f(m_{t,n})|\leq h(m_{t,n},\mathcal{O}_{t}).$
Many contextual bandit algorithms like Lin-UCB (Chu et al., 2011), UCB-GLM (Li et al., 2017), IGP-UCB (Chowdhury and Gopalan, 2017), GP-TS (Chowdhury and Gopalan, 2017), Neural-UCB (Zhou et al., 2020), and Neural-TS (Zhang et al., 2021) are OFD compatible. The value of $h(m_{t,n},\mathcal{O}_{t})$ provides an upper bound on the goodness function of the estimated utility with respect to the true utility function. This value depends on the problem and the choice of contextual bandit algorithm $\mathfrak{A}$ and its associated hyperparameters (e.g., $\delta$; see Table 1 in the Appendix for more details). For a given problem, any appropriate OFD compatible contextual bandit algorithm can be used as a sub-routine to get optimistic utility estimates for all item-agent pairs. These estimates are used to compute the goodness function value, which is used to select the best agent for allocating the given item.
Let $\mathfrak{A}$ be an OFD contextual bandit algorithm with $|f_{t}^{\mathfrak{A}}(m_{t,n})-f(m_{t,n})|\leq h(m_{t,n},\mathcal{O}_{t})$. Then, the agent for allocating the item $m_{t}$ is selected by maximizing the following goodness function: $n_{t}\in\arg\!\max_{n\in\mathcal{N}}\textrm{G}\left(\bm{U}_{t,n}^{\mathfrak{A}}\right),$ where $\bm{U}_{t,n}^{\mathfrak{A}}=\left(U_{t}^{a}+u_{m_{t},a}^{\mathfrak{A}}\mathbbmss{1}{\left(a=n\right)}\right)_{a\in\mathcal{N}}$ in which $U_{t}^{a}$ is the total utility collected by agent $a$ thus far and $u_{m_{t},a}^{\mathfrak{A}}$ is the optimistic estimate of $f(m_{t,a})$ (for example, $u_{m_{t},a}^{\mathfrak{A}}=f_{t}^{\mathfrak{A}}(m_{t,a})+h(m_{t,a},\mathcal{O}_{t})$ for UCB-based algorithms). Note that the assumptions underlying contextual bandit algorithms need to satisfy in our setting, as they directly influence the performance of our proposed algorithm via $h(m_{t,n_{t}},\mathcal{O}_{t})$. Next, we will give a regret upper bound for using any OFD compatible contextual bandit algorithm $\mathfrak{A}$ to get an optimistic utility estimate.
Theorem 2.
Let $\mathfrak{A}$ be an OFD compatible contextual bandit algorithm with $|f_{t}^{\mathfrak{A}}(m_{t,n})-f(m_{t,n})|\leq h(m_{t,n},\mathcal{O}_{t})$ and the goodness function be same as defined in Eq. 2 with $w_{\max}=\max_{n\in\mathcal{N}}{w_{n}}$. If assumptions underlying $\mathfrak{A}$ holds, then, with a probability of at least $1-\delta$, the regret of corresponding OFD algorithm OFD-$\mathfrak{A}$ in $T$ rounds is
$\mathfrak{R}_{T}\left(\textnormal{{OFD-$\mathfrak{A}$}}\right)\leq 2w_{\max}\sqrt{T}\sqrt{\sum_{t=1}^{T}\left[h(m_{t,n_{t}},\mathcal{O}_{t})\right]^{2}}.$
Proof outline.
The proof follows by upper-bounding the sum of instantaneous regrets. Specifically, we first show that instantaneous regret $(r_{t})$ incurred in the round $t$ is upper bounded by $2w_{n_{t}}h(m_{t,n_{t}},\mathcal{O}_{t})$ (as shown in Appendix C), using the fact that the estimation error is $|f_{t}^{\mathfrak{A}}(m_{t,n})-f(m_{t,n})|\leq h(m_{t,n},\mathcal{O}_{t})$ for an item-agent pair $m_{t,n}$ when using an OFD compatible contextual bandit algorithm $\mathfrak{A}$. Thus, the regret of OFD-$\mathfrak{A}$ directly depends on $\mathfrak{A}$ through $h(m_{t,n_{t}},\mathcal{O}_{t})$, resulting in the same order of regret as that of $\mathfrak{A}$, since $w_{\max}$ is a scale-free constant independent of $N$.
4.3 Locally Monotonically Non-decreasing and Lipschitz Goodness Functions
The regret upper bounds are given in 1 and 2 hold for a specific choice of goodness function (defined in Eq. 2). We have also derived the regret upper bounds for goodness functions that satisfy the properties of locally monotonically non-decreasing and locally Lipschitz continuous.
Definition 2 (Locally monotonically non-decreasing and Lipschitz continuous function).
Let $\bm{U}_{t,-n}$ denote the utility of all agents except agent $n$. A goodness function G is
(i) locally monotonically non-decreasing for any $u>0\in\mathbb{R}$ if
$\textrm{G}\left(U_{t,n},\bm{U}_{t,-n}\right)\leq\textrm{G}\left(U_{t,n}+u,\bm{U}_{t,-n}\right),\text{ and}$
(ii) locally Lipschitz continuous if, for a constant $c_{n}>0$ corresponding to the $n$-th element of $\bm{U}$,
$|\textrm{G}\left(U_{t,n},\bm{U}_{t,-n}\right)-\textrm{G}\left(U^{\prime}_{t,n},\bm{U}_{t,-n}\right)|\leq c_{n}|U_{t,n}-U^{\prime}_{t,n}|.$
Next, we present an upper bound on the regret for goodness functions that are locally monotonically non-decreasing and locally Lipschitz continuous, such as NSW and log-NSW.
Theorem 3.
Let $\mathfrak{A}$ be an OFD compatible contextual bandit algorithm with $|f_{t}^{\mathfrak{A}}(m_{t,n})-f(m_{t,n})|\leq h(m_{t,n},\mathcal{O}_{t})$ and the goodness function G be is locally monotonically non-decreasing and locally Lipschitz continuous, with $c_{\max}=\max_{n\in\mathcal{N}}{c_{n}}$. If assumptions used in $\mathfrak{A}$ holds, then, with a probability of at least $1-\delta$, the regret of corresponding OFD algorithm OFD-$\mathfrak{A}$ in $T$ rounds is
$\mathfrak{R}_{T}\left(\textnormal{{OFD-$\mathfrak{A}$}, {G}}\right)\leq 2c_{\max}\sqrt{T}\sqrt{\sum_{t=1}^{T}\left[h(m_{t,n_{t}},\mathcal{O}_{t})\right]^{2}}.$
Proof outline.
The instantaneous regret $(r_{t})$ is upper bounded by $2c_{n_{t}}h(m_{t,n_{t}},\mathcal{O}_{t})$ (as shown in Appendix C). This result uses the following three facts: (i) the locally monotonically non-decreasing property, (ii) the locally Lipschitz continuous property of the goodness function, and (iii) the estimation error bound $|f_{t}^{\mathfrak{A}}(m_{t,n})-f(m_{t,n})|\leq h(m_{t,n},\mathcal{O}_{t})$ for the contextual bandit algorithm $\mathfrak{A}$.
5 Experiments
In this section, we aim to corroborate our theoretical results and empirically demonstrate the performance of our proposed algorithms in different online fair allocation problems. We repeat all our experiments 20 times and show the regret (as defined in Eq. 1) with a 95% confidence interval (the vertical line on each curve shows the confidence interval). To demonstrate the different performance aspects of our proposed algorithms, we have used different synthetic problem instances (commonly used experiment choices in bandit literature) whose details are as follows.
Experiment setting.  We use a $d_{m}$-dimensional space to generate the sample features of each item, where item $m_{t}$ is represented by $m_{t}=\left(x_{m_{t},1},\ldots,x_{m_{t},d_{m}}\right)$ for $t\geq 1$. Similarly, we use a $d_{n}$-dimensional space to generate the sample features of each agent, where agent $n\in\mathcal{N}$ is represented by $n=\left(x_{n,1},\ldots,x_{n,d_{n}}\right)$ represent the agent $n$. The value of $i$-the feature $x_{m_{t},i}$ (or $x_{n,i})$ is sampled uniformly at random from $\left(0,10\right)$. Note that agents remain the same across the rounds, whereas an item in each round is randomly sampled from the $d_{m}$-dimensional space. To get the item-agent feature vectors for item $m_{t}$ in the round $t$, we concatenate the item features $m_{t}$ with all agent feature vectors. For item $m_{t}$ and agent $n$, the concatenated feature vector is denoted by $m_{t,n}$, which is an $d$-dimensional vector with $d=d_{m}+d_{n}$. We select a $d$-dimensional vector $\theta^{\star}$ by sampling uniformly at random from $(0,10)^{d}$ and normalizing it to have unit $l_{2}$-norm. In all experiments, we use $\lambda=0.01$, $R=0.1$, $\delta=0.05$, and $d_{m}=d_{n}$.
Regret comparison with baselines.  To the best of our knowledge, this paper is the first work to model online fair division of numerous items with few copies using a contextual bandit framework. We compare the regret of proposed algorithms with two other baselines: OFD-Uniform and OFD-Greedy. OFD-Uniform selects an agent uniformly at random to allocate the item. In contrast, OFD-Greedy uses $\varepsilon$-Greedy contextual algorithm, which behaves the same as OFD-UCB except it has no confidence term (i.e., setting $\alpha_{t}=0$) and selects an agent uniformly at random with probability $0.1$ in each round, otherwise select agent greedily. For experiments with the linear utility (i.e., $f(x)=x^{\top}\theta^{\star}$), we use $10000$ items, $10$ agents, and $\rho=0.85$ control parameter for the goodness function. We use three different problems with the same setting except $d_{m}=d_{n}=\{2,5,10\}$, resulting in $d=\{4,10,20\}$. As expected, our algorithms based on UCB and TS-based contextual linear bandit algorithms outperform both baselines as shown in Fig. 2a-2c on different problem instances of linear utility (only varying the dimension $d$ while keeping remaining parameters unchanged). Note that we set a limit on the y-axis to highlight the sub-linear regret of our algorithm. Further, we observe that the TS-based algorithm performs better than the UCB-based algorithm.
(a) Linear $(d=4)$
(b) Linear $(d=10)$
(c) Linear $(d=20)$
(d) Varying copies (UCB)
(e) $\min\limits_{n\in\mathcal{N}}U_{t}^{n}/\sum_{n\in\mathcal{N}}U_{t}^{n}$
(f) Gini coefficient
(g) Total utility
(h) Varying copies (TS)
Figure 2: (2a-2c): Comparing cumulative regret of our proposed online fair division algorithms with baseline algorithms. (2d & 2h): Cumulative regret of OFD-UCB and OFD-TS vs. different number of copies for each item $(c)$. (2f-2g): Fairness and efficiency measures (ratio of minimum utility to total utility, Gini coefficient, and total utility) as a function of control parameter $(\rho)$.
Regret vs. number of item’s copies.  Multiple copies of each item make estimating the utility function easier, leading to lower regret. To verify this, we examine how the regret of our proposed algorithms changes as the number of copies per item ($c$) increases. In our experiments, we use a linear utility function ($f(x)=x^{\top}\theta^{\star}$), a total of $1000/c$ items, $N=10$, $\rho=0.85$, $d=40$, and vary $c$ from the set $\{10,25,50,100,200\}$. As expected, the regret of our UCB- and TS-based algorithms decreases with more copies per item, as shown in Fig. 2d and Fig. 2h.
[…]
Regret vs. number of agents ($N$) and dimension $(d)$.  The number of agents $(N)$ and dimension of item-agent feature vector $(d)$ in the online fair division problem control the difficulty. As their values increase, the problem becomes more difficult, making it harder to allocate the item to the best agent We want to verify this by observing how the regret of our proposed algorithms changes while varying $N$ and $d$ in the online fair division problem. To see this in our experiments, we use the linear utility function (i.e., $f(x)=x^{\top}\theta^{\star}$), $1000$ items, $N=10$ when varying dimension, $\rho=1$, $d=40$ while varying the number of agents. As shown in Fig. 3a and Fig. 3b, the regret bound of our UCB- and TS- based algorithms increases as we increase the number of agents, i.e., $N=\{5,10,15,20,25\}$. We also observe the same trend when increasing the dimension of the item-agent feature vector from $d=\{10,20,30,40,50\}$ as shown in Fig. 3c and Fig. 3d. In all experiments, we also observe that the TS-based algorithm performs better than its UCB-based counterpart (as seen in Fig. 3a-3d by comparing the regret of both algorithms). Additional results are provided in Appendix D.
(a) Vary agents (UCB)
(b) Vary agents (TS)
(c) Vary dimension (UCB)
(d) Vary dimension (TS)
Figure 3: Cumulative regret of OFD-UCB and OFD-TS vs. different values of $N$ and $d$.
6 Conclusion
This paper considers a novel variant of the online fair division problem in which there are a large number of items with only a few copies for each items. These items arrive sequentially, and a learner must irrevocably allocate each indivisible item to one of the agents while satisfying predefined fairness and efficiency constraints. Since many real-life applications involve a large number of items with only a few copies of each, it becomes difficult to accurately estimate the utility for all item-agent pairs. To address this challenge, we assume utility is an unknown function of item-agent features. We then propose algorithms with sub-linear regret guarantees that model online fair division as a contextual bandit problem to get optimistic utility estimates for each item-agent pair and use them for item allocation to agents. Our experimental results further validate the performance of the proposed algorithms. For future work, it would be interesting direction to explore alternative fairness notions, such as envy-freeness and proportionality. Another promising direction is to study fairness across multiple types of items in the online fair division setting with unknown utility functions.
Appendix A Additional Related Work
Due to space constraints in the main paper, we will now highlight relevant work in fair division and fair multi-armed bandits pertaining to our problem setting.
Fair division.  The fair division is a fundamental problem (Steinhaus, 1948; Dubins and Spanier, 1961) in which a designer allocates resources to agents with diverse preferences. There are two popular notions of ‘fairness’ used in fair division: envy-freeness (EF) (Varian, 1974) (every agent prefers his or her allocation most) and proportional fairness (Steinhaus, 1948) (each agent is guaranteed her proportional share in terms of total utility independent of other agents). Many classical studies on fair division focus on divisible resources (Steinhaus, 1948; Dubins and Spanier, 1961; Varian, 1974), which are later extended to indivisible items (Brams and Taylor, 1996; Moulin, 2004; Budish, 2011) due to their many practical applications. However, obtaining envy-free allocation may not be guaranteed to exist, e.g., in the case of two agents and one item. Therefore, the appropriate relaxations of EF and proportionality, such as envy-freeness up to one good (EF1) (Moulin, 2004), envy-freeness up to any good (EFX)  (Caragiannis et al., 2019), and maximin share fairness (MMS) (Budish, 2011), are used in practice. We refer the readers to Amanatidis et al. (2022) for a survey on fair division.
Fair multi-armed bandits (MAB).  The fair MAB is a framework where an arm with a higher expected reward is selected with a lower probability than an arm with a lower expected reward (Joseph et al., 2016). The fair policy should sample arms with probability proportional to the value of a merit function of its mean reward (Wang and Joachims, 2021). Many works also assume that preserving fairness means the probability of selecting each arm should be similar if the two arms have a similar quality distribution (Liu et al., 2017; Chen et al., 2021). The fairness is also considered in the linear contextual bandits setting (Gillen et al., 2018; Wu et al., 2023; Schumann et al., 2019; Wang et al., 2021; Grazzi et al., 2022) where an unknown similarity metric imposes individual fairness constraints. Some fair MAB variant seeks to optimize the cumulative reward while also ensuring that, at any round, each arm is pulled at least a specified fraction of times (Li et al., 2019; Chen et al., 2020; Claure et al., 2020; Patil et al., 2021). The work of Hossain et al. (2021) considers a multi-agent setting: the goal is to find a fair distribution over the arms for the agents. Recently, Barman et al. (2023) study Nash regret in the context of stochastic multi-armed bandits with bounded rewards and provide optimal regret guarantees, while Sawarni et al. (2023) extends this analysis to linear bandits.
Appendix B Discussion about Regret defined in Eq. 1
Goodness function.  Since the utility function $f$ (hence $G(\cdot)$) is unknown, our proposed algorithms need to estimate the agent’s utility to compute the goodness function’s value before allocating a given item. For this, we can use a suitable contextual bandit algorithm that gives us an optimistic utility estimate for each item-agent pair. The value of goodness function $G$ for allocating the item $m_{t}$ to an agent $n$ is denoted by $G(U_{t,n})$, where $U_{t,n}$ denotes the vector of all agents’ total utility at the beginning of round $t$, except the total utility of n-th agent is $U_{t}^{n}+f(m_{t},n)$, i.e., optimistic estimate of utility if item $m_{t}$ will be allocated to an agent $n$.
Regret definition.  Our goal is fair allocation each round, motivated by practical applications where learners ensure fairness in item distribution (Procaccia et al., 2024; Sim et al., 2021). For example, on an online platform, service providers (agents) may leave or switch to competitors if they receive too few users. Thus, our algorithms optimize the goodness function each round to balance fairness and efficiency. By design, this definition of regret satisfies the Markovian property; that is, the item allocation to an agent depends only on the current total utility (sum of the observed noisy utilities, i.e., $y_{t}$’s) of the agents and an optimistic estimate of their utility.
Furthermore, the oracle uses the function $f$ to compute the goodness function $G$ for each agent before allocating an item to the best agent. Since $y_{t}$ represents the observed noisy utility after allocation in round $t$, the oracle does not know the observation noise $\varepsilon_{t}$. As a result, Oracle can not use $U_{t}^{n}+f(m_{t},n)+\varepsilon_{t}$ to compute the goodness function $G$ for $n$-th agent. Furthermore, it is a standard practice to ignore noise $\varepsilon_{t}$ when computing the optimal arm in contextual bandit algorithms.
Appendix C Regret Analysis.
We first state the following result that we will use in the proof of 1.
Lemma 1 (Theorem 2 of Abbasi-Yadkori et al. (2011)).
Let $\delta\in(0,1)$, $\lambda>0$, $\hat{\theta}_{t}={M}_{t}^{-1}\sum_{s=1}^{t}m_{s,n}y_{n_{s}}$, $\left\|\theta^{\star}\right\|_{2}\leq S$, and $\left\|m_{t,n}\right\|_{2}\leq L~{}\forall t\geq 1,n\in\mathcal{N}$. Then, with probability $1-\delta$,
$\left\|\hat{\theta}_{t}-\theta^{\star}\right\|_{M_{t}}\leq\left(R\sqrt{d\log\left(\frac{1+\left({tL^{2}}/{\lambda}\right)}{\delta}\right)}+\lambda^{\frac{1}{2}}S\right)=\alpha_{t}.$
We next state the following result that we will use to prove our regret upper bounds for 1 and 2.
Lemma 2 (Lemma 1 of Weymark (1981)).
Let $\bm{w}=(w_{1},w_{2},\ldots,w_{N})$ and $\bm{u}=(u_{1},u_{2},\ldots,u_{N})$ are opposite ordered. If $\bm{U}=\{\bm{u}^{1},\bm{u}^{2},\ldots,\bm{u}^{p}\}$ is the set of all permutations of $\bm{u}$, then
$\sum_{n=1}^{N}w_{n}u_{n}\leq\sum_{n=1}^{N}w_{n}u_{n}^{j},~{}~{}\textnormal{ for }j=1,2,\ldots,p.$
Lemma 3 (Lemma 1 of Sim et al. (2021)).
Consider a non-decreasing order of weights $w_{1}>\ldots>w_{N}$ and let $\textrm{G}\left(\bm{U}\right)=\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}\right)$, where $\bm{U}$ is an $N$-dimensional utility vector and $\Phi_{n}$ returns the $n$-th smallest element of $\bm{U}$. For any two utility vectors $\bm{U}^{a}$ and $\bm{U}^{b}$, if there exists $i$ such that $U_{i}^{a}>U_{i}^{b}$ and $\forall n\in\mathcal{N}\setminus\{i\}$, $U_{n}^{a}=U_{n}^{b}$, then $\textrm{G}\left(\bm{U}^{b}\right)>\textrm{G}\left(\bm{U}^{b}\right)$.
The regret analysis of our proposed algorithms depends on bounding the instantaneous regret for each action. The following result gives an upper bound on the instantaneous regret when using an contextual algorithm $\mathfrak{A}$.
Lemma 4.
Let G be the goodness function defined in Eq. 2 and $\mathfrak{A}$ be an OFD compatible contextual bandit algorithm with $|f_{t}^{\mathfrak{A}}(m_{t,n})-f(m_{t,n})|\leq h(m_{t,n},\mathcal{O}_{t})$. Then, the instantaneous regret incurred by UCB-$\mathfrak{A}$ for selecting agent $n_{t}$ for item $m_{t}$ is
$r_{t}=\textrm{G}\left(\bm{U}_{t,n_{t}^{\star}}\right)-\textrm{G}\left(\bm{U}_{t,n_{t}}\right)\leq 2w_{n_{t}}h(m_{t,n_{t}},\mathcal{O}_{t}).$
Proof.
Recall the definition of the goodness function, i.e., $\textrm{G}\left(\bm{U}_{t,n_{t}}\right)=\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n_{t}}\right)$.
$r_{t}(\mathfrak{A})$ $=\textrm{G}\left(\bm{U}_{t,n_{t}^{\star}}\right)-\textrm{G}\left(\bm{U}_{t,n_{t}}\right)$
$=\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n_{t}^{\star}}\right)-\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n_{t}}\right)$
$=\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}^{\star}\left(\bm{U}_{t,n}\right)+w_{n_{t}^{\star}}f(m_{t,n_{t}^{\star}})-\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n_{t}}\right)$
$\leq\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}^{\star}\left(\bm{U}_{t,n}\right)+w_{n_{t}^{\star}}\left[f_{t}^{\mathfrak{A}}(m_{t,n_{t}^{\star}})+h(m_{t,n_{t}^{\star}},\mathcal{O}_{t})\right]-\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n_{t}}\right)$
$\leq\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n}+\left[f_{t}^{\mathfrak{A}}(m_{t,n_{t}^{\star}})+h(m_{t,n_{t}^{\star}},\mathcal{O}_{t})\right]\mathbbmss{1}{\left(n=n_{t}^{\star}\right)}\right)-\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n_{t}}\right)$
$\leq\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n}+\left[f_{t}^{\mathfrak{A}}(m_{t,n_{t}})+h(m_{t,n_{t}},\mathcal{O}_{t})\right]\mathbbmss{1}{\left(n=n_{t}\right)}\right)-\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n_{t}}\right)$
$\leq\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}^{f_{t}}\left(\bm{U}_{t,n}+\left[f_{t}^{\mathfrak{A}}(m_{t,n_{t}})+h(m_{t,n_{t}},\mathcal{O}_{t})\right]\mathbbmss{1}{\left(n=n_{t}\right)}\right)-\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n_{t}}\right)$
$=\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}^{f_{t}}\left(\bm{U}_{t,n}\right)+w_{n_{t}}\left[f_{t}^{\mathfrak{A}}(m_{t,n_{t}})+h(m_{t,n_{t}},\mathcal{O}_{t})\right]-\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n_{t}}\right)$
[…]
$=w_{n_{t}}\left[f_{t}^{\mathfrak{A}}(m_{t,n_{t}})+h(m_{t,n_{t}},\mathcal{O}_{t})\right]-w_{n_{t}}f(m_{t,n_{t}})$
$=w_{n_{t}}\left[f_{t}^{\mathfrak{A}}(m_{t,n_{t}})-f(m_{t,n_{t}})+h(m_{t,n_{t}},\mathcal{O}_{t})\right]$
$\leq w_{n_{t}}\left[|f_{t}^{\mathfrak{A}}(m_{t,n_{t}})-f(m_{t,n_{t}})|+h(m_{t,n_{t}},\mathcal{O}_{t})\right]$
$\leq w_{n_{t}}\left[h(m_{t,n_{t}},\mathcal{O}_{t})+h(m_{t,n_{t}},\mathcal{O}_{t})\right]$
$=2w_{n_{t}}h(m_{t,n_{t}},\mathcal{O}_{t}).$
The $\Phi_{n}^{\star}(\cdot)$ returns the utilities in the same order as $\Phi_{n}\left(\bm{U}_{t,n_{t}^{\star}}\right)$, while $\Phi_{n}^{f_{t}}(\cdot)$ returns the utilities in the same order as $\Phi_{n}\left(\bm{U}_{t,n_{t}}\right)$. The first inequality is from applying $|f_{t}^{\mathfrak{A}}(m_{t,n})-f(m_{t,n})|\leq h(m_{t,n},\mathcal{O}_{t})$ with $m_{t,n}=m_{t,n_{t}^{\star}}$ and then using triangle inequality to get $f(m_{t,n_{t}^{\star}})\leq f_{t}^{\mathfrak{A}}(m_{t,n_{t}^{\star}})+h(m_{t,n_{t}^{\star}},\mathcal{O}_{t})$. The second inequality follows from 3 as an additional utility of $f_{t}^{\mathfrak{A}}(m_{t,n_{t}^{\star}})+h(m_{t,n_{t}^{\star}},\mathcal{O}_{t})$ is added to the total collected utility of agent $n_{t}^{\star}$. The third inequality follows from the fact that $n_{t}$ maximizes the goodness function. The fourth inequality follows from 2. The last inequality is due to $|f_{t}^{\mathfrak{A}}(m_{t,n})-f(m_{t,n})|\leq h(m_{t,n},\mathcal{O}_{t})$. This completes the proof, showing that $r_{t}(\mathfrak{A})\leq 2w_{n_{t}}h(m_{t,n_{t}},\mathcal{O}_{t})$. ∎
C.1 Proof of 1
See 1
Proof.
As $h(m_{t,n_{t}},\mathcal{O}_{t})=\left(R\sqrt{d\log\left(\frac{1+\left({tL^{2}}/{\lambda}\right)}{\delta}\right)}+\lambda^{\frac{1}{2}}S\right)\left\|m_{t,n_{t}}\right\|_{{M}_{t}^{-1}}$ when linear contextual bandit algorithm is used. After using 2, we have
[…]
$=2w_{\max}\sqrt{T}\sqrt{\sum_{t=1}^{T}\left[\left(R\sqrt{d\log\left(\frac{1+\left({tL^{2}}/{\lambda}\right)}{\delta}\right)}+\lambda^{\frac{1}{2}}S\right)\left\|m_{t,n_{t}}\right\|_{{M}_{t}^{-1}}\right]^{2}}$
$=2w_{\max}\sqrt{T}\sqrt{\sum_{t=1}^{T}\left[\left(R\sqrt{d\log\left(\frac{1+\left({tL^{2}}/{\lambda}\right)}{\delta}\right)}+\lambda^{\frac{1}{2}}S\right)\left\|m_{t,n_{t}}\right\|_{{M}_{t}^{-1}}\right]^{2}}$
$=2w_{\max}\sqrt{T}\sqrt{\sum_{t=1}^{T}\left(R\sqrt{d\log\left(\frac{1+\left({tL^{2}}/{\lambda}\right)}{\delta}\right)}+\lambda^{\frac{1}{2}}S\right)^{2}\left[\left\|m_{t,n_{t}}\right\|_{{M}_{t}^{-1}}\right]^{2}}$
$=2w_{\max}\sqrt{T}\left(R\sqrt{d\log\left(\frac{1+\left({tL^{2}}/{\lambda}\right)}{\delta}\right)}+\lambda^{\frac{1}{2}}S\right)\sqrt{\sum_{t=1}^{T}\left[\left\|m_{t,n_{t}}\right\|_{{M}_{t}^{-1}}\right]^{2}}$
$\leq 2w_{\max}\sqrt{T}\left(R\sqrt{d\log\left(\frac{1+\left({TL^{2}}/{\lambda}\right)}{\delta}\right)}+\lambda^{\frac{1}{2}}S\right)\sqrt{\sum_{t=1}^{T}\left[\left\|m_{t,n_{t}}\right\|_{{M}_{t}^{-1}}\right]^{2}}$
$=2\alpha_{T}w_{\max}\sqrt{T}\sqrt{\sum_{t=1}^{T}\left[\left\|m_{t,n_{t}}\right\|_{{M}_{t}^{-1}}\right]^{2}}$
$\leq 2\alpha_{T}w_{\max}\sqrt{T}\sqrt{2\log\frac{\textnormal{det}(M_{T})}{\textnormal{det}(\lambda I_{d})}}$
$\leq 2\alpha_{T}w_{\max}\sqrt{2dT\log(\lambda+TL/d)}$
[…]
The last two inequalities follow from Lemma 11 and Lemma 10 of Abbasi-Yadkori et al. (2011), respectively. With this, the proof of 1 is complete. ∎
C.2 Proof of 2
See 2
Proof.
Let $r_{t}(\mathfrak{A})$ denote the instantaneous regret for using contextual bandit algorithm $\mathfrak{A}$. After using 4, the regret with contextual bandit algorithm $\mathfrak{A}$ after $T$ rounds is given as follows:
$\mathfrak{R}_{T}(\mathfrak{A})$ $=\sum_{t=1}^{T}r_{t}(\mathfrak{A})\leq 2\sum_{t=1}^{T}w_{n_{t}}h(m_{t,n_{t}},\mathcal{O}_{t})\leq 2w_{\max}\sum_{t=1}^{T}h(m_{t,n_{t}},\mathcal{O}_{t})$
$\leq 2w_{\max}\sqrt{T\sum_{t=1}^{T}\left[h(m_{t,n_{t}},\mathcal{O}_{t})\right]^{2}}$
$\implies\mathfrak{R}_{T}(\mathfrak{A})$ $\leq 2w_{\max}\sqrt{T}\sqrt{\sum_{t=1}^{T}\left[h(m_{t,n_{t}},\mathcal{O}_{t})\right]^{2}}.\qed$
C.3 Proof of 3
See 3
Proof.
First we first get upper bound on instantaneous regret:
$r_{t}(\mathfrak{A})$ $=\textrm{G}\left(\bm{U}_{t,n_{t}^{\star}}\right)-\textrm{G}\left(\bm{U}_{t,n_{t}}\right)$
$=\textrm{G}\left(U_{t,n_{1}},\ldots,U_{t,n_{t}^{\star}}+f(m_{t,n_{t}^{\star}}),\ldots,U_{t,n_{N}}\right)-\textrm{G}\left(\bm{U}_{t,n_{t}}\right)$
$\leq\textrm{G}\left(U_{t,n_{1}},\ldots,U_{t,n_{t}^{\star}}+f_{t}^{\mathfrak{A}}(m_{t,n_{t}^{\star}})+h(m_{t,n_{t}^{\star}},\mathcal{O}_{t}),\ldots,U_{t,n_{N}}\right)-\textrm{G}\left(\bm{U}_{t,n_{t}}\right)$
$\leq\textrm{G}\left(U_{t,n_{1}},\ldots,U_{t,n_{t}}+f_{t}^{\mathfrak{A}}(m_{t,n_{t}})+h(m_{t,n_{t}},\mathcal{O}_{t}),\ldots,U_{t,n_{N}}\right)-\textrm{G}\left(\bm{U}_{t,n_{t}}\right)$
$=\textrm{G}\left(U_{t,n_{1}},\ldots,U_{t,n_{t}}+f_{t}^{\mathfrak{A}}(m_{t,n_{t}})+h(m_{t,n_{t}},\mathcal{O}_{t}),\ldots,U_{t,n_{N}}\right)$
$\qquad\qquad-\textrm{G}\left(U_{t,n_{1}},\ldots,U_{t,n_{t}}+f(m_{t,n}),\ldots,U_{t,n_{N}}\right)$
[…]
$=c_{n_{t}}\left[f_{t}^{\mathfrak{A}}(m_{t,n_{t}})-f(m_{t,n_{t}})+h(m_{t,n_{t}},\mathcal{O}_{t})\right]$
$\leq c_{n_{t}}\left[|f_{t}^{\mathfrak{A}}(m_{t,n_{t}})-f(m_{t,n_{t}})|+h(m_{t,n_{t}},\mathcal{O}_{t})\right]$
$\leq c_{n_{t}}\left[h(m_{t,n_{t}},\mathcal{O}_{t})+h(m_{t,n_{t}},\mathcal{O}_{t})\right]=2c_{n_{t}}h(m_{t,n_{t}},\mathcal{O}_{t}).$
The first inequality follows from the monotonicity property of goodness function after applying $|f_{t}^{\mathfrak{A}}(m_{t,n})-f(m_{t,n})|\leq h(m_{t,n},\mathcal{O}_{t})$ with $m_{t,n}=m_{t,n_{t}^{\star}}$ and then using triangle inequality to get $f(m_{t,n_{t}^{\star}})\leq f_{t}^{\mathfrak{A}}(m_{t,n_{t}^{\star}})+h(m_{t,n_{t}^{\star}},\mathcal{O}_{t})$. The second inequality follows from the fact that $n_{t}$ maximizes the goodness function. The third inequality is due to the locally Lipschitz property of goodness function. Let $c_{\max}=\max\{c_{1},c_{2},\ldots,c_{N}\}$. Then, after $T$ rounds, the regret with contextual bandit algorithm $\mathfrak{A}$ is given as follows:
$\mathfrak{R}_{T}(\mathfrak{A})$ $=\sum_{t=1}^{T}r_{t}(\mathfrak{A})$
$\leq 2\sum_{t=1}^{T}c_{n_{t}}h(m_{t,n_{t}},\mathcal{O}_{t})$
$\leq 2c_{\max}\sqrt{T\sum_{t=1}^{T}\left[h(m_{t,n_{t}},\mathcal{O}_{t})\right]^{2}}$
$\implies\mathfrak{R}_{T}(\mathfrak{A})$ $\leq 2c_{\max}\sqrt{T}\sqrt{\sum_{t=1}^{T}\left[h(m_{t,n_{t}},\mathcal{O}_{t})\right]^{2}}.\qed$
Table 1: Examples of different $h(x,\mathcal{O}_{t})$ values for some contextual bandit algorithms (using notations from original papers).
Contextual bandit algorithm $h(x,\mathcal{O}_{t})$
Lin-UCB (Chu et al., 2011) $\left(R\sqrt{d\log\left(\frac{1+\frac{tL^{2}}{\lambda}}{\delta}\right)}+\lambda^{\frac{1}{2}}S\right)\left\|x\right\|_{{M}_{t}^{-1}}$
GLM-UCB (Li et al., 2017) $\sqrt{\frac{d}{2}\log(1+2t/d)+\log(1/\delta)}\frac{\left\|x\right\|_{{M}_{t}^{-1}}}{\kappa}$
IGP-UCB (Chowdhury and Gopalan, 2017) $\sqrt{2(\gamma_{t-1}+1+\log(1/\delta))}\sigma_{t-1}(x)$ + $B\sigma_{t-1}(x)$
Appendix D Additional Experimental Results
The number of agents $(N)$ and dimension of item-agent feature vector $(d)$ in the online fair division problem control the difficulty. As their values increase, the problem becomes more difficult, making it harder to allocate the item to the best agent. We want to verify this by observing how the regret of our proposed algorithms changes while changing $N$ and $d$ in the online fair division problem. To see this in our experiments, we use the linear utility function (i.e., $f(x)=x^{\top}\theta^{\star}$), $1000$ items, $N=10$ when varying dimension, $\rho=1$, $d=40$ while varying the number of agents. As shown in Fig. 4a and Fig. 4b, the regret bound of our UCB- and TS- based algorithms increases as we increase the number of agents, i.e., $N=\{5,10,15,20,25\}$. We also observe the same trend when we increase the dimension of the item-agent feature vector from $d=\{10,20,30,40,50\}$ as shown in Fig. 4c and Fig. 4d. In all experiments, we also observe that TS-based algorithm performs better than its UCB-based counterpart (as seen in Fig. 4 by comparing the regret of both algorithms). However, when we set the value of $\rho=0.85$, we observe the same trend for $d$, but the trend reverses as the number of agents increases due to the goodness function (defined in Eq. 2) decreases with an increase in the number of agents when $\rho<1$, as shown in Fig. 3. It happens because the value of the goodness function decreases with an increase in the number of agents when $\rho<1$.
(a) Vary agents (UCB)
(b) Vary agents (TS)
(c) Vary dimension (UCB)
(d) Vary dimension (TS)
Figure 4: Cumulative regret of OFD-UCB and OFD-TS vs. different values of $N$ and $d$ for $\rho=1.0$.
Non-linear utility function.  For this experiment, we adapt problem instances with non-linear utility functions from those used for linear utility functions in Section 5. We apply a polynomial kernel of degree $2$ to transform the item-agent feature vectors to introduce non-linearity. The constant terms (i.e., the $1$’s) resulting from this transformation are removed. As an example, a sample $4$-d feature vector $x=(x_{1},x_{2},x_{3},x_{4})$ is transformed into a $14$-d feature vector: $x^{\prime}=(x_{1},x_{2},x_{3},x_{4},x_{1}x_{2},x_{1}x_{3},x_{1}x_{4},x_{2}x_{3},x_{2}x_{4},x_{3}x_{4},x_{1}x_{2}x_{3},x_{1}x_{2}x_{4},x_{1}x_{3}x_{4},x_{2}x_{3}x_{4})$. We also remove $1$’s, which appears in the transformed samples. As shown in Fig. 5, our algorithms (nOFD-UCB and nOFD-TS, prefixed with ‘n’) consistently outperform both baselines across various non-linear problem instances (where only the feature dimension $d=\{4,6,10,20\}$ is varied, while all other parameters remain same as the instances for linear utility functions). To better demonstrate the sub-linear regret behavior of our algorithms, we limit the y-axis in the plots. Additionally, we observe that the TS-based algorithm achieves lower regret than its UCB-based counterpart.
(a) $d=4$
(b) $d=6$
(c) $d=10$
(d) $d=20$
Figure 5: Cumulative regret of OFD-UCB and OFD-TS vs. different values $d$ for $\rho=0.85$.
Computational resources.  All the experiments are run on a server with AMD EPYC 7543 32-Core Processor, 256GB RAM, and 8 GeForce RTX 3080.
Appendix E Auxiliary Observations
We first mention some of the observations that will be useful to prove the property of monotonicity and Lipschitz continuity for our various choices of the goodness function.
Fact 1.
Suppose $h(x)$ and $g(x)$ are Lipschitz continuous functions with non-negative bounded domains. Then the following statements are true:
  1. 1.
    $h(x)+g(x)$ is a Lipschitz continuous function.
  2. 2.
    $h(x)g(x)$ is a Lipschitz continuous function.
Proof.
Let $h(x)$ and $g(x)$ be the Lipschitz continuous functions in a bounded domain with Lipschitz constants $k_{1}$ and $k_{2}$, respectively.
1. We first prove that $h(x)+g(x)$ is a Lipschitz continuous function. Consider the following difference:
$\bigg{|}(h(x)+g(x))-(h(y)+g(y))\bigg{|}$ $\leq\bigg{|}h(x)-h(y)\bigg{|}+\bigg{|}g(x)-g(y)\bigg{|}$
$\leq k_{1}\bigg{|}x-y\bigg{|}+k_{2}\bigg{|}x-y\bigg{|}$
$\leq k\bigg{|}x-y\bigg{|}$
where $k=k_{1}+k_{2}$. The first inequality follows from the triangle inequality, while the second inequality uses Lipschitz continuity.
2. We now prove that $h(x)g(x)$ is a Lipschitz continuous function. Consider the following difference:
$\bigg{|}h(x)g(x)-h(y)g(y)\bigg{|}$ $=$ $\bigg{|}h(x)g(x)-h(y)g(x)+h(y)g(x)-h(y)g(y)\bigg{|}$
$\leq$ $\bigg{|}h(x)g(x)-h(y)g(x)\bigg{|}+\bigg{|}h(y)g(x)-h(y)g(y)\bigg{|}$
$\leq$ $|g(x)|\bigg{|}h(x)-h(y)\bigg{|}+|h(y)|\bigg{|}g(x)-g(y)\bigg{|}$
$\leq$ $Mk_{1}\bigg{|}x-y\bigg{|}+Mk_{2}\bigg{|}x-y\bigg{|}\leq M^{\prime}\bigg{|}x-y\bigg{|}~{}\mbox{ where }M^{\prime}=Mk.$
The last inequality is due to the bound of the continuous function, and appropriately, the Lipschitz constant is chosen. ∎
Fact 2.
Suppose $h_{i}(x)$ with $1\leq i\leq n$ are Lipschitz continuous functions with non-negative bounded domain. Then the following statements are true:
  1. 1.
    $\sum_{i=1}^{n}a_{i}h_{i}(x)$ is a Lipschitz continuous function, where all $a_{i}^{\prime}s$ are constant.
  2. 2.
    $\prod_{i=1}^{n}h_{i}(x)$ is a Lipschitz continuous function.
Proof.
1. We first prove that $\sum_{i=1}^{n}a_{i}h_{i}(x)$ is a Lipschitz continuous function. Consider the following difference:
$\bigg{|}\sum_{i=1}^{n}a_{i}h_{i}(x)-\sum_{i=1}^{n}a_{i}h_{i}(y)\bigg{|}$ $=$ $\bigg{|}\sum_{i=1}^{n}a_{i}\bigg{(}h_{i}(x)-h_{i}(y)\bigg{)}\bigg{|}$
$\leq$ $\sum_{i=1}^{n}a_{i}\bigg{|}h_{i}(x)-h_{i}(y)\bigg{|}$
$\leq$ $\sum_{i=1}^{n}k_{i}a_{i}\bigg{|}x-y\bigg{|}=L\bigg{|}x-y\bigg{|}.$
The first inequality is due to the triangle inequality, while the second inequality is by using Lipschitz continuity of the function.
2. We now prove that $\prod_{i=1}^{n}h_{i}(x)$ is a Lipschitz continuous function. We will use induction for this proof. For base case: When $n=2$, the statement is true due to Fact 1. For induction hypothesis: Assume the statement is true for $n=m$. Now, we consider the inductive step (for $n=m+1$) as follows:
$\bigg{|}\prod_{i=1}^{m+1}h_{i}(x)-\prod_{i=1}^{m+1}h_{i}(y)\bigg{|}$
$\qquad=\bigg{|}\prod_{i=1}^{m+1}h_{i}(x)-h_{m+1}(y)\prod_{i=1}^{m}h_{i}(x)+h_{m+1}(y)\prod_{i=1}^{m}h_{i}(x)-\prod_{i=1}^{m+1}h_{i}(y)\bigg{|}$
$\qquad=\bigg{|}\bigg{(}h_{m+1}(x)-h_{m+1}(y)\bigg{)}\prod_{i=1}^{m}h_{i}(x)+h_{m+1}(y)\bigg{(}\prod_{i=1}^{m}h_{i}(x)-\prod_{i=1}^{m}h_{i}(y)\bigg{)}\bigg{|}$
$\qquad\leq\bigg{|}\bigg{(}h_{m+1(x})-h_{m+1}(y)\bigg{)}\prod_{i=1}^{m}h_{i}(x)\bigg{|}+\bigg{|}h_{m+1}(y)\bigg{(}\prod_{i=1}^{m}h_{i}(x)-\prod_{i=1}^{m}h_{i}(y)\bigg{)}\bigg{|}$
$\qquad=\bigg{|}\prod_{i=1}^{m}h_{i}(x)\bigg{|}\bigg{|}\bigg{(}h_{m+1}(x)-h_{m+1}(y)\bigg{)}\bigg{|}+\bigg{|}h_{m+1}(y)\bigg{|}\bigg{|}\bigg{(}\prod_{i=1}^{m}h_{i}(x)-\prod_{i=1}^{m}h_{i}(y)\bigg{)}\bigg{|}$
$\qquad\leq\beta\bigg{|}x-y\bigg{|}.$
The first inequality follows from the triangle inequality. The last inequality is due to the boundedness of the function combined with the induction hypothesis. ∎
Fact 3.
The function $log(x)$ is Lipschitz continuous when $x$ is in a non-negative domain, i.e., $x>0$.
Proof.
Suppose $0<x\leq l$. Now consider the following difference.
$\bigg{|}\log(x)-log(y)\bigg{|}$ $=\bigg{|}\log\bigg{(}\frac{y}{x}\bigg{)}\bigg{|}=\bigg{|}\log\bigg{(}1+\frac{y}{x}-1\bigg{)}\bigg{|}\leq\bigg{|}\frac{y}{x}-1\bigg{|}\leq\frac{1}{l}\bigg{|}x-y\bigg{|}.$
We use that $z>\log(1+z)$ for $z>-1$. Therefore, it proves our claim. ∎
Fact 4.
Suppose $h(x)$ is a continuous function, bounded and positive, then $log(h(x))$ is a Lipschitz continuous function.
Proof.
Suppose $0<h(x)\leq l$. Now consider the following difference.
$\bigg{|}\log\bigg{(}h(x)\bigg{)}-\log\bigg{(}h(y)\bigg{)}\bigg{|}$ $=\bigg{|}\log\bigg{(}\frac{h(y)}{h(x)}\bigg{)}\bigg{|}$
$=\bigg{|}\log\bigg{(}1+\frac{h(y)}{h(x)}-1\bigg{)}\bigg{|}$
$\leq\bigg{|}\frac{h(y)}{h(x)}-1\bigg{|}\leq\frac{1}{l}\bigg{|}h(x)-h(y)\bigg{|}.$
Using Fact 3, the proof is complete. ∎
E.1 Goodness functions that are locally monotonically non-decreasing and Lipschitz continuous
We will briefly discuss the different choices of the goodness function for which we can theoretically guarantee a sub-linear regret upper bound.
Weighted Gini social-evaluation function.  The weighted Gini social-evaluation function is given as follows:
$\textrm{G}\left(\bm{U}_{t,n_{t}}\right)=\sum_{n\in\mathcal{N}}w_{n}~{}\Phi_{n}\left(\bm{U}_{t,n_{t}}\right).$
Since all $w_{n}$’s are positive constants with $w_{\max}=\max_{n\in N}w_{n}$ and $\Phi_{n}\left(\bm{U}_{t,n_{t}}\right)$ are bounded, positive, continuous functions. Therefore, the weighted Gini social-evaluation function satisfies the locally Lipschitz condition. Also, if we change the $i$-th component while keeping the rest of the component of the weighted Gini social-evaluation function fixed, the locally monotonicity non-decreasing condition holds (also from 3). Therefore, 3 (with $c_{\max}=w_{\max}$) also holds for weighted Gini social-evaluation function.
Targeted weights.  Let us define a fixed target weight vector in advance, with the learner’s goal being to achieve these targeted weight distributions after each allocation. These target weights represent the desired fraction of the total cumulative utility each agent should receive. Let $\text{SU}_{n_{t}}=f(m_{t},n_{t})+\sum_{n\in N}U_{t}^{n}$ be the sum of total utility after allocating item $m_{t}$ to agent $n_{t}$. Suppose the target weighted fraction of the utility vector is ${\bf r^{*}}$, which is given. The learner’s goal is to obtain the targeted weight vector at the end of the allocation process. The goodness function for this fairness constraint is defined as follows:
$\textrm{G}\left({\bf U}_{t,n_{t}}\right)=\sum_{n\in N}w_{n}~{}\Phi_{n}({\bf U}_{t,n_{t}}^{\bf p}),$
where $w_{1}=1$ and $w_{i}=0$ for $i\geq 2$ [i.e., ESW]. The utility ${U}_{t,n}^{\bf p}$ in ${\bf U}_{t,n_{t}}^{\bf p}$ is defined as ${U}_{t,n}^{\bf p}={U}_{t,n}/p_{n}$, where $p_{n}$ is the proportional ratio for the $n$-th agent. Let ${\bf r^{*}}=(r_{1},\ldots,r_{N})$ be the vector of targeted ratio of agents’ cumulative utility to total utilities collected by all agents (i.e., system’s total utility), where $\sum_{n=1}^{N}r_{n}=1$. Then, $p_{n}=r_{n}/\min_{i}(r_{i})$, e.g., if $N=3$ and ${\bf r}=(0.2,0.5,0.3)$, then ${\bf p}=(0.2/0.2,0.5/0.2,0.3/0.2)=(1,2.5,1.5)$.
Nash Social Welfare.  The Nash Social Welfare (NSW) is defined as the product of the utilities of all agents, i.e., $\left[\prod_{n\in\mathcal{N}}U_{n}\right]$. Since we are maximizing the goodness function, we can ignore the constant factor $\frac{1}{|\mathcal{N}|}$ here. First, observe that the utility function is $\textrm{G}(\cdot):\mathbb{R}^{n}\mapsto\mathbb{R}^{+}$. Since the utility function is positive, NSW satisfies the monotonically non-decreasing function property. Also, we assume that the utility of the individual agent is positive and bounded, which implies that the NSW is locally Lipschitz continuous. As a result, the properties given in 2 hold for NSW.
Log Nash Social Welfare.  This function is defined as the $\log$ of the Nash social welfare function and is commonly considered in the fair division literature (Cole and Gkatzelis, 2018; Talebi and Proutiere, 2018).
$\textrm{G}\left(\bm{U}_{t,n_{t}}\right)$ $=\log\left[\prod_{n\in\mathcal{N}}\left(U_{t}^{n}+f(m_{t},n_{t})\mathbbmss{1}{\left(n=n_{t}\right)}\right)\right]=\sum_{n\in\mathcal{N}}\log(U_{t}^{n}+f(m_{t},n_{t})\mathbbmss{1}{\left(n=n_{t}\right)}).$
We assume that the utility function is bounded and positive. It is easy to observe that using Fact 1- Fact 4, the $\log$ NSW is a Lipschitz continuous function. Also, $\log$ NSW satisfies the monotone property. By virtue of these observations, we can guarantee sub-linear regret, i.e., 3 holds.